DYNAMICS OF ACETYLENE ISOTOPOMERS REVEALED BY FREQUENCY ANALYSIS

M. JACOBSON, Department of Chemistry, Columbia University, New York, NY 10027; Z. DUAN, M. SILVA and \underline{R. W. FIELD}, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

Analysis of the Dispersed Fluorescence spectra of Acetylene Isotopomers (^{12}C_{2}H_{2}, ^{13}C_{2}H_{2}, ^{12}C_{2}HD) has continued to provide information about the structure and dynamics on both the X̃ state and Ã state potential surfaces. Sampling of specific regions of the potential surfaces is a consequence of the different integer multiple ratios among the fundamental vibrational frequencies of the different isotopomers. Insights from these studies are summarized.